The current price of a non-dividend-paying stock is 30. The volatility of the stock is 0.3 per annum. The risk free rate is 0.05 for all maturities. Using the Cox-Ross-Rubinstein binomial tree model with two time steps to do the valuation, what is the value of a European call option with a strike price of 32 that expires in 6 months? (Your answer should be in the unit of dollar (up to the precision of cents), but without the dollar sign. For example, if your answer is $1.02, just enter 1.02. Please also think whether your answer will be different if the option is an American option.)
Answer #1

Details provided : risk free rate = 5%, current price = 30, volatility = 30%, 6 month European call option = 32, t = 3/12 - since this is a two step binomial tree

Solution:

u = e^(volatility*squareroot(time)) = 1.16183

d = 1/u = 0.86

a = e^rt = 1.012578

p = (a-d)/(u-d) = 0.504343

Call price = 3.315